OpenStudy (anonymous):

help please.... find the values of x where the extreme values of the function y=x^3-7x-5x occur.

OpenStudy (anonymous):

that not my answer choice

OpenStudy (anonymous):

(a) x=-0.33, x=5 (b) x=0.33, x=-5 (c) x=0.33, x=5 (d) x=2, x=5 (e) x=0.65 ,x=7.65

OpenStudy (nali):

I think you forgot to put x^2 somewhere in ur question. Edit your question and plug in ur new formula and check if is is right.

OpenStudy (nali):

If your question is y=x^3-7x^2-5x, then your right choice might be a and check from the website @nihalpetal gave you how to do the problem

OpenStudy (nali):

@nihalpatel